OpenStudy (anonymous):

can you please give the equation for this problem? The numerator of a rational number exceeds it denominator by 3. If 1 is added to the denominator and 5 is subtracted from the numerator, the value of the rational number becomes 1/2. What is the rational number?

OpenStudy (anonymous):

suppose denominator of the rational no is x then fraction= (x+3)/x (x+3-5)/(x+1)= 1/2

Directrix (directrix):

The numerator of a rational number exceeds it denominator by 3. --- (x + 3)/x If 1 is added to the denominator and 5 is subtracted from the numerator --- (x + 3 -5) / (x + 1) = 1/2 (x-2)/(x + 1) = 1/2 x = 5 so the original rational number is (5+3)/5 = 8/5.

Directrix (directrix):

@danielle02 Question? The task was to find the rational number.

OpenStudy (anonymous):

@Directrix yes, the rational number is being asked

Directrix (directrix):

@danielle02 Do you agree that the rational number = 8/5 ?

OpenStudy (anonymous):

@Directrix the rational number is 5. not 8/5

Directrix (directrix):

The solution to this equation (x-2)/(x + 1) = 1/2 is x = 5 but x is not the original rational number. The original rational number is described here: The numerator of a rational number exceeds it denominator by 3" It would look like (x + 3 ) / x. x turned out to be 5. Substituting, (5 + 3) /x = 8/5 for the rational number. ------ Check it. 8/5 If 1 is added to the denominator and 5 is subtracted from the numerator -->rational number becomes 1/2 (8 - 5) / (5 + 1) = 3/6 = 1/2. 8/5 works.

Directrix (directrix):

@danielle02

OpenStudy (anonymous):

thanks for your help :) i checked it again and it turns out, you were right.

Directrix (directrix):

Okay. Have you done the age problem yet?